zunairah:

help with math x+y=(-4) y=4x+1

boredfr:

Ok so do you know how to start this off?

zunairah:

no

boredfr:

Ok so we are gonna solve it using substitution Rewrite your equation: \[y=4x+1;x+y=4\]

zunairah:

ok

boredfr:

Then you are going to solve for y: \[Substitute:4x+1\] for y in \[x+y=4\]

boredfr:

\[x+y=4\]\[x+4x+1=4\]\[5x+1=4\]\[5x+1+1=4+1\]\[5x=-5\]\[\frac{5x }{ 5 }=\frac{ -5 }{ 5 }\] Which means x equals what ?

zunairah:

1

zunairah:

i mean -1

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zunairah 1 \(\color{#0cbb34}{\text{End of Quote}}\) Correct now you have to solve for y aswell correct?

zunairah:

yes

boredfr:

Ok so now that we have x= -1 we will start on y Substitute -1 for x in y=4x+1 \[y=4x+1\]\[y=(4)(-1)+1\] Now if you simplify both sides of the equations then y should equal?

zunairah:

-3

zunairah:

thx